OpenStudy (anonymous):

I have some coins, and 15 of them are nickels. The ratio of nickels to dimes is 3 : 2, and the ratio of dimes to quarters is 5 : 4. How many quarters do I have?

OpenStudy (texaschic101):

3/2 = 15/x (x being dimes) cross multiply 3x = 30 x = 10 (dimes) 5/4 = 10/x (x being quarters) cross multiply 5x = 40 x = 8 (quarters) you have 8 quarters

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

Great job, that maes so much sense

OpenStudy (anonymous):

:D